OpenStudy (anonymous):

The equation y = 1777x + 27,153 can be used to predict the number y of gun deaths in the US x years after 2000, that is, x = 0 corresponds to 2000, x = 3 corresponds to 2003, x = 5 corresponds to 2005, and son on. Predict the number of gun deaths in 2005 and 2009. In what year will the number of gun deaths be 9383? Gun deaths in 2005? Gun deaths in 2009? Gun deaths will be 9383 in the year??

OpenStudy (anonymous):

Subtract 2000 from each of those years and plug that value in for X. then plug and chug out the math to get Y for each one.

OpenStudy (anonymous):

36,038 43,146 x=-10 so 1990

OpenStudy (saifoo.khan):

hewweo, im back! xD

OpenStudy (anonymous):

Good, Nice to have you back

OpenStudy (anonymous):

Doing my final

OpenStudy (saifoo.khan):

need help?

OpenStudy (anonymous):

yea.....

OpenStudy (saifoo.khan):

Alright, so let's start.

OpenStudy (anonymous):

ty

OpenStudy (saifoo.khan):

uw

OpenStudy (anonymous):

The qeustion I have posted right now is a big one.

OpenStudy (saifoo.khan):

this one right?

OpenStudy (anonymous):

right.

OpenStudy (saifoo.khan):

in 2005, 5 years = y = 1777x + 27,153 y = 1777(5) + 27,153 y = 36 038

OpenStudy (saifoo.khan):

In 2009, 9 years = y = 1777x + 27,153 y = 1777(9) + 27,153 y = 43 146

OpenStudy (saifoo.khan):

In last part, solve for x, 9383 = 1777x + 27,153

OpenStudy (anonymous):

Doesn't make sense. With thos calcuations, the number of gun deaths will go down rather than up.

OpenStudy (saifoo.khan):

LOL. why they will come down?

OpenStudy (anonymous):

36,000 in 2005....43,000 in 2009, and 9,383 in the year???

OpenStudy (anonymous):

The equation in every day terms says that gun deaths increase by 1777 per year. With gun deaths being at 27,153 in the year 2000.

OpenStudy (saifoo.khan):

I doubt that the equation u typed in is incorrect.

OpenStudy (saifoo.khan):

y = 1777x + 27,153 y = -1777x + 27,153 which one?

OpenStudy (anonymous):

Maybe I typed it wrong

OpenStudy (saifoo.khan):

LOL, so which one now?

OpenStudy (anonymous):

Well if it's negative then that makes more sense. Because when it's positive you'll get 1990 as the year with 9383 deaths. Which is invalid assuming the trend (equation) only applys to 2000 onwards. The year will be 2010 if it's -1777

OpenStudy (anonymous):

Let me do this again.... y = -1777x + 27153 to predict the # of gun deaths in the US x yrs after 2000, that is, x=0 corresponds to 2000, x=3 corresponds to 2003, x = 5 corresponds to 2005, and so on. Predict the # of gun deaths in 2005 and 2009. In what year will the # of gun deaths be 9383?

OpenStudy (saifoo.khan):

in 2005, 5 years = y = -1777x + 27,153 y = -1777(5) + 27,153 y = 18 268 In 2009, 9 years = y = -1777x + 27,153 y = -1777(9) + 27,153 y = 11 160

OpenStudy (saifoo.khan):

In the last one, y = -1777x + 27,153 9383 = -1777x + 27,153 x = 10 means after 10 years. 2000+10 = 2010.

OpenStudy (anonymous):

So the # of gun deaths have gone down?

OpenStudy (saifoo.khan):

Right, now.

OpenStudy (anonymous):

Well, I only have 2 qustions left and have 10 minutes to get it turned in, so I think I am just going to forget those last two

OpenStudy (saifoo.khan):

c'mon post them, i will give u the answers right away!

OpenStudy (anonymous):

One is one we have done before

OpenStudy (saifoo.khan):

which one? and second?

OpenStudy (anonymous):

soybean meal is 16 percent protein and cornmeal is 8 percent proteien. How many lbs of each should be mixed togegether in order to get 320 lbs mixure that is 15% protein?

OpenStudy (saifoo.khan):

280 = the amount of soybean meal 40 = cornmeal.

OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (4,3) and (9,-2)

OpenStudy (saifoo.khan):

\[\frac{-2-3}{9-4}\]\[\frac{-5}{5}\]\[-1\]

OpenStudy (saifoo.khan):

Please hit good answer ^^^

OpenStudy (saifoo.khan):

So how much did u scored?

OpenStudy (anonymous):

79%

OpenStudy (anonymous):

Not as good as I wanted, but that's okay.

OpenStudy (anonymous):

I would have totally bombed had it not been for your help.

OpenStudy (saifoo.khan):

=(

OpenStudy (anonymous):

Tomorrow I start Algebra 1B.....CRUD!!!!

OpenStudy (saifoo.khan):

Oh, cool. i will be around for help.

OpenStudy (anonymous):

Cool? You are insane!!! LOL

OpenStudy (saifoo.khan):

HAHA, why is that?

OpenStudy (anonymous):

Anyone who thinks that math is "cool" is insane!!! hahaha

OpenStudy (saifoo.khan):

HAHA.

OpenStudy (anonymous):

well, gotta get to bed. It's3:18 am and I have to get my son off to school in about 4 hours...arrgghhh. Thanks for you help.